Session: Sep 1999 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (mauro at Thu Sep 16 02:01:58 METDST 1999) Success
All scans is recorded. gps-formatter= -44.8 uSec the sky clear

Torun (K.Borkowski at Thu Sep 16 09:35:53 METDST 1999) Success
Till 21:00 UT no calibration signal.

Effelsberg (A. Kraus at Thu Sep 16 10:08:42 METDST 1999) Minor failures
The first two scans were lost due to a breakdown of the elevation drive.

Westerbork (Tony Foley at Thu Sep 16 11:02:01 METDST 1999) Minor failures
Scans until 18:00UT lost due to hardware problem in tied-array A/D converters. Thunder and lightning reset some telescopes during the run and one telescope (RTB) stopped moving.

Noto (C.Trigilio at Thu Sep 16 17:40:38 METDST 1999) Minor failures
The recorder lost vacuum several times. This problem occurred at the begin of passes 1,2,4,5,6,7,8,10 and in those cases the tape has been started manually. A new tape was mounted before pass 5, but the problem remained. Tapes sent on 16 Sep morning.

Medicina ( at Fri Sep 17 10:45:16 METDST 1999) Minor failures

Shanghai (Huang Xinyong at Fri Sep 17 10:50:21 METDST 1999) Success
all scans were recorded. clock offset=UTC(sheshan25)-UTC(gps)=1.98us. It is cloudy.

Jodrell (Alastair Gunn at Mon Sep 20 11:08:15 METDST 1999) Minor failures
Telescope control system failed at approximately 19:05 UT. System was driven manually to each source after coming back online at about 21:24 UT. Some data may also be missing after this time due to ongoing problems with telescope control.

Calibration (Jean\-Francois Desmurs at Sun Nov 7 15:38:43 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the sep99 directory under the name: c99c3.antab. Read the comments in the file to get further information.